The FBI wants to determine the effectiveness of their 10 Most Wanted list. To do so, they need to find out the fraction of people who appear on the list that are actually caught. Suppose a sample of 977 suspected criminals is drawn. Of these people, 400 were captured. Using the data, estimate the proportion of people who were caught after being on the 10 Most Wanted list.

  • A. 0.040

  • B. 0.100

  • C. 0.977

  • D. 0.400

  • E. 0.409
Answer #1

Solution :

Given that,

x = 400

n = 977

Point estimate = sample proportion = \hat p = x / n = 400 / 977 = 0.409

correct option is = E